OpenStudy (anonymous):

Solve for x: 12x +16y = 6 a. x=1/2-3/4y b. x= 1/2-4/3y c. x= 2-3/4y d. x= 2+4/3y

OpenStudy (texaschic101):

12x + 16y = 6 --- subtract 16y from both sides 12x = -16y + 6 --- divide both sides by 12 x = (-16/12)y + (6/12) x = -4/3y + 1/2......rearranged to match answer choice : x = 1/2 - 4/3y any questions ?

OpenStudy (texaschic101):

basically, you have to get x by itself. And remember, what you do on one side of the equal sign, you have to do on the other.

OpenStudy (anonymous):

Alright Thank you